OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2+8x-4=0

OpenStudy (anonymous):

do you know how to "complete a square"? first add \(4\) to both sides and start with \[x^2+8x=4\] now what is half of \(8\)?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok good, so we write \[(x+4)^2=4+?\]now what is \(4^2\) ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

then that is what goes on the right side \[(x+4)^2=4+16\\ (x+4)^2=20\]

OpenStudy (anonymous):

now take the square root, don't forget the \(\pm\) and write \[x+4=\pm\sqrt{20}\\x+4=\pm2\sqrt5\]

OpenStudy (anonymous):

and finally subtract \(4\) from both sides and get \[x=-4\pm2\sqrt5\] now it is done

OpenStudy (anonymous):

ThankYou satellite73 !